OpenStudy (buttermequeasy):

Find the Mistake: Provided below is a solution to the equation 8x + 7 − 2x = 27 − 2. Choose the statement that is true about the given solution. Line # 8x + 7 − 2x = 27 − 2 1 6x + 7 = 25 2 x + 7 = 19 3 x = 12

OpenStudy (buttermequeasy):

The provided solution is wrong. There is a mistake in line #1. The provided solution is wrong. There is a mistake in line #2. The provided solution is wrong. There is a mistake in line #3. The entire solution is correct.

OpenStudy (saifoo.khan):

it's in front!!! SPOT IT.

OpenStudy (buttermequeasy):

Aha I see it.

OpenStudy (anonymous):

Line # 8x + 7 − 2x = 27 − 2 1 6x + 7 = 25 (Should not work on both sides of equation) 2 x + 7 = 19 (Must Subtract Before Dividing and Cannot subtract 6 from 25) 3 x = 12 (Answer is wrong because of bad math practices) How it should be done: 1 6x + 7 = 27 - 2 2 6x = 18 3 x = 3